# Model Inversion Attacks on Vision-Language Models: Do They Leak What They Learn?

### v1

@@ 4.2 Initial Candidate Selection
Following the method from (Struppek et al. 2022), we perform an initial selection to identify promising candidates for inversion. We begin by sampling 2000 latent vectors, denoted as $\{w\}_{i=1}^{2}000$, from the prior distribution. For each $w$, we evaluate the target VLMs loss. We then select the top $n$ vectors with the lowest loss to serve as our initialization candidates. In our experiments, we set $n=16$ to create 16 candidates for attacks.
@@ 4.3 Final Selection
To select the final reconstructed image, we perform a final selection step, also following the method from (Struppek et al. 2022). This step aims to identify the reconstructed images that have the highest confidence. For each of the $n$ initialization candidates, we apply 10 random data augmentations and re-evaluate the target VLMs loss. We calculate the average loss for each candidate across these augmentations and select the $n/2$ candidates with the lowest average loss as the final attack outputs.
@@ 5 Related Work
Model Inversion. Model Inversion (MI) seeks to recover information about a model’s private training data via pretrained model. Given a target model $M$ trained on a private dataset $\mathcal{D}_{\text{priv}}$, the adversary aims to infer sensitive information about the data in $\mathcal{D}_{\text{priv}}$, despite it being inaccessible after training. MI attacks are commonly framed as the task of reconstructing an input $x^{r}_{y}$ that the model $M$ would classify as belonging to a particular label $y$. The foundational MI method is introduced in (Fredrikson et al. 2014), demonstrating that machine learning models could be exploited to recover patients’ genomic and demographic data.
@@
Building on the foundational work of (Fredrikson et al. 2014), early MI attacks targeting facial recognition are proposed in (Fredrikson, Jha, and Ristenpart 2015; Yang et al. 2019), demonstrating the feasibility of reconstructing recognizable facial images from the outputs of pretrained models. However, performing direct optimization in the high-dimensional image space is challenging due to the large search space. To address this, recent advanced generative-based MI attacks have shifted the search to the latent space of deep generative models (Zhang et al. 2020; Wang et al. 2021; Chen et al. 2021; Yang et al. 2019; Yuan et al. 2023; Nguyen et al. 2023b; Struppek et al. 2022; Qiu et al. 2024).
@@
Specifically, GMI (Zhang et al. 2020) and PPA (Struppek et al. 2022) employ WGAN (Arjovsky, Chintala, and Bottou 2017) and StyleGAN (Karras, Laine, and Aila 2019), respectively, trained on an auxiliary public dataset $\mathcal{D}_{\text{pub}}$ that similar to the private dataset $\mathcal{D}_{\text{priv}}$. The pretrained GAN is served as prior knowledge for the inversion process. To improve this prior knowledge, KEDMI (Chen et al. 2021) trains inversion-specific GANs using knowledge extracted from the target model $M$. PLGMI (Yuan et al. 2023) introduces pseudo-labels to enhance conditional GAN training. IF-GMI (Qiu et al. 2024) utilizes intermediate feature representations from pretrained GAN blocks. Most recently, PPDG-MI (Peng et al. 2024) improves the generative prior by fine-tuning GANs on high-quality pseudo-private data, thereby increasing the likelihood of sampling reconstructions close to true private data. Beyond improving GAN-based priors, several studies focus on improving the MI objective including max-margin loss (Yuan et al. 2023) and logit loss (Nguyen et al. 2023b) to better guide the inversion process. Additionally, LOMMA (Nguyen et al. 2023b) introduces the concept of augmented models to improve the generalizability of MI attacks.
@@
Unlike MI attacks, MI defenses aim to reduce the leakage of private training data while maintaining strong predictive performance. Several approaches have been proposed to defend against MI attacks. MID (Wang, Zhang, and Jia 2021) and BiDO (Peng et al. 2022) introduce regularization-based defenses that include the term of regularization in the training objective. The crucial drawback of these approaches is that the regularizers often conflict with the training objective resulting in a significant degradation in model’s utility. Beyond regularization-based strategies, TL-DMI (Ho et al. 2024) leverages transfer learning to improve MI robustness, and LS (Struppek, Hintersdorf, and Kersting 2024) applies Negative Label Smoothing to mitigate inversion risks. Architectural approaches to improve MI robustness have also been explored in (Koh et al. 2024). More recently, Trap-MID (Liu and Chen 2024) introduces a novel defense by embedding trapdoor signals into $M$. These signals act as decoys that mislead MI attacks into reconstructing trapdoor triggers instead of actual private data.
@@
Model Inversion in Multimodal Large Vision-Language Models. Large Vision-Language Models (VLMs) are increasingly deployed in many real-world applications across diverse domains, including sensitive areas. Unlike unimodal vision models, VLMs are designed to process both image and text inputs and generate text responses. A typical VLM architecture includes a text tokenizer to encode textual inputs into text tokens, a vision encoder to extract image features as image tokens, and a lightweight projection layer that maps image tokens into the text token space. These tokens are then concatenated and passed through a LLM to produce the final response. This multimodal processing pipeline fundamentally distinguishes VLMs from traditional unimodal vision models.
@@ 6.2 Limitations
While following conventional MI attacks to focus on facial images and dog breeds, a more diverse domain scenarios, such as natural scenes or medical images, remain an important direction for future research. Moreover, evaluations on a broader range of models are needed to further comprehend our study on MI for VLMs.

### v2

@@ 4.3 Initial Candidate Selection
Following the method from [struppek2022plug], we perform an initial selection to identify promising candidates for inversion. We begin by sampling 2000 latent vectors, denoted as $\{w\}_{i=1}^{2000}$, from the prior distribution. For each $w$, we evaluate the target VLMs loss. We then select the top $n$ vectors with the lowest loss to serve as our initialization candidates. In our experiments, we set $n=16$ to create 16 candidates for attacks.
@@ 4.4 Final Selection
To select the final reconstructed image, we perform a final selection step, also following the method from [struppek2022plug]. This step aims to identify the reconstructed images that have the highest confidence. For each of the $n$ initialization candidates, we apply 10 random data augmentations and re-evaluate the target VLMs loss. We calculate the average loss for each candidate across these augmentations and select the $n/2$ candidates with the lowest average loss as the final attack outputs.
@@ 5 Related Work
Model Inversion. Model Inversion (MI) seeks to recover information about a model’s private training data via pretrained model. Given a target model $M$ trained on a private dataset $\mathcal{D}_{\text{priv}}$, the adversary aims to infer sensitive information about the data in $\mathcal{D}_{\text{priv}}$, despite it being inaccessible after training. MI attacks are commonly framed as the task of reconstructing an input that the model $M$ would classify as belonging to a particular label $y$. The foundational MI method is introduced in [fredrikson2014privacy], demonstrating that machine learning models could be exploited to recover patients’ genomic and demographic data.
@@
Building on the foundational work of [fredrikson2014privacy], early MI attacks targeting facial recognition are proposed in [fredrikson2015model, yang2019neural], demonstrating the feasibility of reconstructing recognizable facial images from the outputs of pretrained models. However, performing direct optimization in the high-dimensional image space is challenging due to the large search space. To address this, recent advanced generative-based MI attacks have shifted the search to the latent space of deep generative models [zhang2020secret, wang2021variational, chen2021knowledge, yang2019neural, yuan2023pseudo, nguyen2023re, struppek2022plug, qiu2024closer].
@@
Specifically, GMI [zhang2020secret] and PPA [struppek2022plug] employ WGAN [arjovsky2017wasserstein] and StyleGAN [karras2019style], respectively, trained on an auxiliary public dataset $\mathcal{D}_{\text{pub}}$ that similar to the private dataset $\mathcal{D}_{\text{priv}}$. The pretrained GAN is served as prior knowledge for the inversion process. To improve this prior knowledge, KEDMI [chen2021knowledge] trains inversion-specific GANs using knowledge extracted from the target model $M$. PLGMI [yuan2023pseudo] introduces pseudo-labels to enhance conditional GAN training. IF-GMI [qiu2024closer] utilizes intermediate feature representations from pretrained GAN blocks. Most recently, PPDG-MI [peng2024pseudo] improves the generative prior by fine-tuning GANs on high-quality pseudo-private data, thereby increasing the likelihood of sampling reconstructions close to true private data. Beyond improving GAN-based priors, several studies focus on improving the MI objective including max-margin loss [yuan2023pseudo] and logit loss [nguyen2023re] to better guide the inversion process. Additionally, LOMMA [nguyen2023re] introduces the concept of augmented models to improve the generalizability of MI attacks.
@@
Unlike MI attacks, MI defenses aim to reduce the leakage of private training data while maintaining strong predictive performance. Several approaches have been proposed to defend against MI attacks. MID [wang2021improving] and BiDO [peng2022bilateral] introduce regularization-based defenses that include the term of regularization in the training objective. The crucial drawback of these approaches is that the regularizers often conflict with the training objective resulting in a significant degradation in model’s utility. Beyond regularization-based strategies, TL-DMI [ho2024model] leverages transfer learning to improve MI robustness, and LS [struppekcareful] applies Negative Label Smoothing to mitigate inversion risks. Architectural approaches to improve MI robustness have also been explored in [koh2024vulnerability]. More recently, Trap-MID [liu2024trap] introduces a novel defense by embedding trapdoor signals into $M$. These signals act as decoys that mislead MI attacks into reconstructing trapdoor triggers instead of actual private data.
@@
Model Inversion in Multimodal Large Vision-Language Models. Large Vision-Language Models (VLMs) are increasingly deployed in many real-world applications across diverse domains, including sensitive areas [liu2024llavanext, Qwen2.5-VL, chen2023minigptv2, chen2024internvl, chen2024expanding]. Unlike unimodal vision models, VLMs are designed to process both image and text inputs and generate text responses. A typical VLM architecture includes a text tokenizer to encode textual inputs into text tokens, a vision encoder to extract image features as image tokens, and a lightweight projection layer that maps image tokens into the text token space. These tokens are then concatenated and passed through a LLM to produce the final response. This multimodal processing pipeline fundamentally distinguishes VLMs from traditional unimodal vision models.
@@ 6.2 Limitations
While following conventional MI attacks to focus on facial images and dog breeds, a more diverse domain scenarios, such as natural scenes or medical images, remain an important direction for future research. Moreover, evaluations with more models can further support our claims.